OpenStudy (anonymous):

will fan and medal find the function rule x-2 -1 0 1 2 y9 4 -1 -6 -11

OpenStudy (kkutie7):

so kinda like a sequence right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@Kkutie7

OpenStudy (anonymous):

@d

OpenStudy (anonymous):

@Outh-1

OpenStudy (anonymous):

@Kkutie7

OpenStudy (kkutie7):

well. x is going up by an addition of one.

OpenStudy (anonymous):

ok

OpenStudy (kkutie7):

y looks like its going down by 5. have you tried calculating a slope? if it is constant then we know the next step

OpenStudy (anonymous):

i dont hae a lotof timei will fan and medal you right now for a direct answer please

OpenStudy (kkutie7):

constant slope is equal to -5/1 by taking the points and using y2-y1/x2-x1

OpenStudy (anonymous):

ok

OpenStudy (kkutie7):

so now use y-y1=m(x-x1) y-9=(-5)(x+2)

OpenStudy (kkutie7):

y=-5x-1